OpenStudy (anonymous):

A wall is to be constructed around a rectangle of 240 square yards with a fence running the length. The construction cost for the fence is $5 per yard and for the wall $8 per yard. a.) Determine the model cost of the total project as a function of the length x of the fence.

OpenStudy (campbell_st):

since area = lw or 240 = lw l = 240/w if x = length then width = 240/x perimeter P = 2x + 2(240/x) P = 2x + 480/x now the problem gets confusing is the wall on all 4 sides and fence of 2